OpenStudy (anonymous):

You need to dilute 2.50 mL of a 8.00 M KOH solution to make a 2.00M KOH solution. What is the volume of diluted solution after you add water?

OpenStudy (abmon98):

Use Dilutions Law: C1*V1=C2*V2 2.50*8/1000=2*V2/1000 We Divide by 1000 to Change from ml to dm^3